Grade 12Algebra

Fill in the blanks to make the following statements true:
17 + ............ = 0

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us assume the missing integer be x,
Then,
= 17 + x = 0
By sending 17 from LHS to RHS it becomes -17,
= x = 0 - 17
= x = - 17
Now substitute the x value in the blank place,
= 17 + (-17) = 0 ... [This equation is in the form of Closure property of Addition]
= 17–17 = 0

Thanks